OpenStudy (anonymous):

Using complete the square or -b/2a , what is the vertex of: 1/2(x^2 + 4x -2) Using complete the square or -b/2a , what is the vertex of: 1/2(x^2 + 4x -2) @Mathematics

hero (hero):

I'd go with x = -b/2a to find the vertex

OpenStudy (anonymous):

I keep getting to completely different answers though. For completing the square I get (-1,-3) and for -b/2a I get (-1,-5/2).

hero (hero):

Don't use complete the square for this. Use -b/2a then check it by graphing

OpenStudy (anonymous):

I checked it by graphing, and the answers are closer to (-1,-3). I need someone to walk me through the steps

OpenStudy (anonymous):

I could tell you everthing that I;ve done if that would help. Maybe you could pick out my mistakes

hero (hero):

I did that using -b/2a

hero (hero):

Use .5(x^2+4x-2) instead of the fraction

OpenStudy (anonymous):

Okay. Can you tell me how you did it? When I use that form, I do get -1 to begin with, but after I replace all of the x's with -1 I get a fraction

OpenStudy (anonymous):

my 2 cents completing the square is to much of a pain in the but to do with fractions.

OpenStudy (anonymous):

Fractions don't bother me. It only bothers me that my check didn't match the first answer that I got

OpenStudy (asnaseer):

you can also use completing the square as follows:\[\frac{1}{2(x^2+4x-2)}=\frac{1}{2((x+2)^2-6)}\]so the verticies are found when the denominator is zero, i.e.:\[2((x+2)^2-6)=0\]\[(x+2)^2=6\]\[x=-2\pm\sqrt{6}\]

OpenStudy (asnaseer):

I mean vertical asymptotes

OpenStudy (anonymous):

I need the vertex to start off with, otherwise the whole equation would be wrong

hero (hero):

Are you suggesting that the way he set up the fraction is correct?

OpenStudy (asnaseer):

have I got the equation wrong?

hero (hero):

I get the feeling that this is another fail to use parentheses when typing linear fractions

OpenStudy (anonymous):

I'll just show my work: Let me scan it really quick

OpenStudy (anonymous):

almost done :D

OpenStudy (asnaseer):

ah - ok @hero, reading your previous answer I see the equation could have been this instead:\[0.5(x^2+4x-2)\]is this correct @Ket-kat?

OpenStudy (anonymous):

How're are you getting there?

hero (hero):

asnaseer, might post steps

OpenStudy (anonymous):

@asnaseer: I don't think it is

hero (hero):

I hate this problem

OpenStudy (anonymous):

sorry xD I was fine until the check didn't match up

hero (hero):

-2,-3 final answer

OpenStudy (asnaseer):

ok, in which case we have:\[0.5(x^2+4x-2)=0.5((x+2)^2-6)=0.5(x+2)^2-3\]which is in standard vertex form giving vertex as (-2,-3)

OpenStudy (anonymous):

where did you get minus 6?

OpenStudy (asnaseer):

\[x^2+4x-2=x^2+4x+4-6\]and\[x^2+4x+4=(x+2)^2\]

OpenStudy (anonymous):

? Wouldn't you have to distribute the 1/2 first and then isolate the 3rd variable? Then you subtract 2? because you add 4 the the left side?

OpenStudy (anonymous):

And I say 2 because you'd have to mult the -4 by the 1/2 first? xO

OpenStudy (asnaseer):

-2 = 4 - 6 so I just replace the "-2" with "4-6" so that I could complete the square with the terms left on the left-hand-side

OpenStudy (anonymous):

I'm still lost as to how you're getting the 6.

OpenStudy (asnaseer):

ok, let me try to explain more clearly

OpenStudy (anonymous):

kay' :)

hero (hero):

I hate to say it, but using -b/2a is easier

OpenStudy (anonymous):

xD It can be.

OpenStudy (asnaseer):

\[x^2+4x-2=x^2+4x+(-2)\]\[=x^2+4x+(4-6)\]\[=x^2+4x+4-6\]\[=(x^2+4x+4)-6\]\[=(x+2)^2-6\]is that clearer?

OpenStudy (anonymous):

so you're answer is -2,-6 ?

OpenStudy (asnaseer):

no, what I just showed was how to factorise x^2+4x-2

OpenStudy (asnaseer):

in your question, you also had a factor of 0.5 which needs to be applied

OpenStudy (anonymous):

My twin sis' said that you're right! She just worked it out for me. YOU'RE AMAZING :D

hero (hero):

We already discussed that the vertex was (-2,-3)

OpenStudy (asnaseer):

:-) thank you, I'm glad I was able to convey the essence of the process to you - these things are not easy to grasp or teach. but well done for picking it up in the end.

OpenStudy (anonymous):

I am actually really good at these questions, but my methods didn't work on this particular equation. Thanks for teaching me a new trick :)

hero (hero):

What new trick was that?

OpenStudy (anonymous):

how to get the right answer...

hero (hero):

I think the only mistake you made was not turning 1/2 into .5

OpenStudy (anonymous):

I don't think that was it. They pretty much ignored the 1/2 until the end?

hero (hero):

All you had to do after that was simply distribute the .5 to get .5x^2+2x-1 Then you could have used -b/2a after that

OpenStudy (anonymous):

that's exactly what I was doing, and I still had the wrong answer

hero (hero):

Are you sure? Using .5 is easier than using 1/2 Trust me

hero (hero):

I'm pretty sure you were using 1/2 and NOT .5 the first time around. Trust me they're the same number but not exactly the same computational wise

OpenStudy (anonymous):

did you look at the paper I attached? Maybe you can see the mistake I made?

hero (hero):

I already told you what your mistake was. Use .5 next time

OpenStudy (anonymous):

My teacher takes of 20 points for not using fractionss... she's chinese

hero (hero):

She can't take points off for that. That's silly

OpenStudy (anonymous):

if o.5 and 1/2 are the same thing then it shouldn't make a difference xD

hero (hero):

I take it that you hate decimals

OpenStudy (anonymous):

She's the head of the math department, I'm pretty sure she can

OpenStudy (anonymous):

I'm a lover not a hater

hero (hero):

She probably can, but it isn't right. Math is already difficult. Getting the right answer should be sufficient enough

OpenStudy (asnaseer):

@Ket-kat - you make me laugh - you have an acute sense of humour - seriously :O) @hero and @Ket-kat - there is no right or wrong method - just one that works best for you for the problem given to you

OpenStudy (anonymous):

Well the answer calls for 1/2 to begin with

hero (hero):

1/2 = .5

hero (hero):

The answer doesn't have .5 in it

OpenStudy (anonymous):

@ Asnaseer : MAHAL KITA ;D Same to you Bre

hero (hero):

or 1/2

OpenStudy (anonymous):

My Textbook proves you otherwise, sorry xo

hero (hero):

The problem has 1/2 in it, not the answer. You must be confuzzled

OpenStudy (asnaseer):

@Ket-kat I take it that means you are from the Philippines?

OpenStudy (anonymous):

no, you have an easy time doing that for yourself. That applies to me 100 fold

OpenStudy (anonymous):

I'm from America. I just happen to be a Pinay ;D

OpenStudy (asnaseer):

I have a friend at work who is from the Philippines which is how I recognised what you said immediately :-)

OpenStudy (asnaseer):

take car @Ket-kat

OpenStudy (asnaseer):

care**

OpenStudy (anonymous):

Take care Asnaseer ;D It was a pleasure. Thank you for everthing

OpenStudy (anonymous):

*ultimate fan*

OpenStudy (anonymous):

Thanks to you too Hero :D

hero (hero):

:D

hero (hero):

You know I was just teasing you

OpenStudy (anonymous):

;D here's my nerdy face I:B (buck-teeth)